LHC Commissioning 2017
 Share
The version of the browser you are using is no longer supported. Please upgrade to a supported browser.Dismiss

 
View only
 
 
Still loading...
ABCDEFGHIJKLMNOPQRSTUVWXYZAA
1
Commissioning plan 2017 shift by shift
2
3
DateShiftTimeStatusActivity
4
29/04/2017
5
M4MKI interlock and synchronization tests with beams to TEDs
6
M4MKI interlock and synchronization tests with beams to TEDs
7
Beam !A4Injection, first turn, RF capture B1&B2 - closed orbit
2 hours to get stored beams in both rings
8
A4BI work: BPM phasing, BCT checks, multi-turn & AC dipole
9
N4BI work: BPM phasing, BCT checks, multi-turn & AC dipole
10
N4New optics, cycle to 6.5TeV, then Q, Q' at injection
11
30/04/2017
12
M4Optics measurements at 450GeV / feedback preparation for R&S / BI checks
13
M4Optics measurements at 450GeV / feedback preparation for R&S / BI checks
14
A4Coarse collimator set-up and loss map in preparation for ramp & ULO check
15
A4Ramp and squeeze to 1m, optics at flat top
16
N4Repeat ramp and squeeze to 1m (coupling and FF)
17
N4Balistic optics - beam did not circulate (will have to thread)
18
01/05/2017Monday
19
M4FIDEL measurements, BPM and BI checks SPS MKP issue
20
M4No beam from SPSSPS MKP issue
21
A4Ramp and squeeze to 1m
22
A4Squeeze to 40 cm in steps, optics measurements
First attemp lost at 40cm RQTF.A67B1
23
N4Squeeze to 40 cm in steps, optics measurements and corrections
24
N4MPS tests of BLM system at injection
25
02/05/2017Tuesday
26
M4MPS tests of BLM system at injection
27
M4MKI8 agk setup, MKI interlock tests (BETS & AKG --> extraction)
28
A4Ballistic optics setup
29
A4MKD risetime
30
N4Full cycle flat with probes, squeeze in one step
31
N4Kick response in TLs
32
03/05/2017Wednesday
33
M4Access
34
M4Access
35
A4ADT setup with multiple probe bunches
36
A4ADT setup with multiple probe bunches
37
N4Ballistic optics, issues to get the machine ready FIDEL issue?
38
N4Ballistic optics
39
04/05/2017Thursday
40
M4Ballistic optics
41
M4Ballistic optics
42
A4Injection of nominal bunch - RF, BI
43
A4ADT setup for TF
44
N4K-modulation of BPMs in triplets
45
N4K-modulation of BPMs in triplets
46
05/05/2017Friday
47
M4RF cavity phasing
as of 10 AM, estiated to 2-3 hours (confirmed by Helga)
48
M4RF cavity phasing
Access needed to UX45
49
A4Optics measurement and correction at 40 & 33 cm
Access for sextupole in S78
50
A4Optics measurement and correction at 40 & 33 cm
51
N4Optics measurement and correction at 40 & 33 cm
52
N4Ramp down issue
53
06/05/2017Saturday
54
M4Reference orbit with nominal bunch
55
M4Ramp and squeeze to 30cm with flat orbit, orbit cleaning
56
A4BTF and BBLRCryo issue Pt6
57
A4MKI waveform
58
N4MKI waveform
59
N4Dump at 2 TeV
60
07/05/2017Sunday
61
M4Reference orbit with nominal bunch and bump
62
M4Aperture at 450 GeV - TBC
63
A4Ramp and squeeze with bumps, aperture at 40 cm
64
A4Ramp and squeeze with bumps, aperture at 40 cm
65
N4Arm & Dump
66
N4Arm & Dump
67
08/05/2017Monday
68
M4Arm & Dump; Problems with various circuits and acrive circuit RQ5.LR3
69
M4Measurements on NLs at 40cm
Trip SIS interlock on RCBX current
70
A4Measurements on NLs at 40cm
Issues 60A powering permit (PIC issue)
71
A4Measurements on NLs at 40cm
72
N4Measurements on NLs at 40cm
73
N4Arm & Dump without beam -- Almost completed
74
09/05/2017Tuesday
75
M2Check steering; BSRT checks at injection
Issue with BPM capture!
76
M2Splashes for ATLAS & CMS & ALICE
77
M4Splashes for ATLAS & CMS & ALICE
78
A4Collimation setup at injection
79
A4Collimation setup at injection
80
N4TDI and TCDQ setup at injection
81
N4Chroma and Arm & Dump without beam -- Completed
82
10/05/2017Wednesday
83
M4Planned access
84
M4Pilot squeeze and collision
85
A2Pilot squeeze and collision
86
A2BPM tests (capture)
87
A4Transfer line collimator setup TI2
88
N2Transfer line collimator setup TI8
89
N2FMCM tests at injection
90
N4Optics measurements/corrections 33cm/30cm or MPS tests (FMCM test in the end)
If beam still available, EOF FMCM test
91
11/05/2017Thursday
92
M4MPS (BPM interlock)
93
M4RF setting up (nominal ramp)
94
A6Nominal squeeze and collisions - 1b per beam: collisions ATLAS/CMS
If beam still available, EOF FMCM test
95
A2Transfer line collimator setup TI8
96
N4Validation TCDIs TI8
97
N4NL corrections and final optics at 40cm
If beam still available, EOF FMCM test
98
12/05/2017Friday
99
M6Aperture measurement of dump and injection region (450GeV)
If beam still available, EOF FMCM test
100
M2Access - EN/CV IP7 (water leak)
Loading...
 
 
 
2017 SHIFT-by-SHIFT - realized
Parameters 2017
2017 Commissioning action
2016 SHIFT-by-SHIFT
2016 Commissioning actions
Collimation
BI